OpenStudy (anonymous):

if f(x) = 1/x how do u find 2f(x)+3 and 3f(x-1)+2 in simplest form?

OpenStudy (anonymous):

well u said f(x) = 1/x so 2f(x) +3 should become 2/x + 3 => (2 + 3x) --------- x now try another one on ur own !

OpenStudy (callisto):

2f(x)+3 = 2(1/x) +3 = (2+3x)/x 3f(x-1)+2 = 3[1/(x-1)]+2 = [3+(2x-2)] / (x-1) = (2x+1) / (x-1)

OpenStudy (anonymous):

very good Jabelle !

OpenStudy (callisto):

acutally, just replace x by the value inside the bracket. eg. when f(x) = 1/x f(x-1) = 1/(x-1) f(x+2) = 1/(x+2) and when you see 2f(x), then put the value of f(x) there ie 2 x (1/x) = 2/x